OpenStudy (anonymous):

each time a touchdown is scored at a football game 6 points are added to the score of the scoring team. a team already has 12 points what rule represents the number of points as an arithmetic sequence when n is the number of touchdowns scored a. A(n)=12+6n b.A(n)12-(n-1)6 c.A(n)=12+(n-1) d.A(n)=12+(n-6)

OpenStudy (maryfitzgerald14):

each time a touchdown is scored at a football game 6 points are added to the score of the scoring team. a team already has 12 points what rule represents the number of points as an arithmetic sequence when n is the number of touchdowns scored a. A(n)=12+6n b.A(n)12-(n-1)6 c.A(n)=12+(n-1) d.A(n)=12+(n-6)